OpenStudy (anonymous):

WHAT IS THE VALUE OF (1/8)^2?

OpenStudy (jonnyvonny):

1/8= .125 (1/8)^2=(.125)^2=.015625

OpenStudy (anonymous):

I NEED THE ANSWER IN FRACTION

OpenStudy (ddcamp):